OpenStudy (anonymous):

Can someone help me with: Five times a number k minus 9 is two thirds of a number j. Solve for j.

OpenStudy (anonymous):

5k - 9 = 2/3 of j. Does this sound about right? Lol

OpenStudy (anonymous):

5k= 2/3J +9

OpenStudy (anonymous):

It also wants it solved.. I got the equation.. It's just the solving

OpenStudy (anonymous):

2 Variables. Any way to make it one makes it easy to solve

OpenStudy (anonymous):

It wants J solved